the distance s along z axis varies with time as a s = 2t 3 +3z 2 +8. find the time t after which acceleration becomes zero
1
Expert's answer
2019-08-01T09:57:28-0400

The law of motion

"s(t)=2t^3+3t^2+8"

The acceleration

"a(t)=\frac{d^2s(t)}{dt^2}=12t+6"

"12t+6=0"

"t=-0.5\:\rm{s}"